OpenStudy (anonymous):

how many multiples of 4 lie between 11 and 266

OpenStudy (anonymous):

266/4=66.5

OpenStudy (anonymous):

so then 11/4=2.75

OpenStudy (anonymous):

so 66.5-2.75=63.75

OpenStudy (anonymous):

or 63 3/4

OpenStudy (anonymous):

It has to be a full number, so isn't it 63 or 64?

OpenStudy (anonymous):

haha oh yeh so its 64 if you want a full number

OpenStudy (anonymous):

Yes, it's 64.

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

wheres my medal?!

OpenStudy (anonymous):

thanks